# Cooperative Edge Caching Based on Elastic Federated and Multi-Agent Deep Reinforcement Learning in Next-Generation Networks

### v1

@@ IV-A2 Local Training
The reconstruction stage which mainly focuses on training the AEN is first executed. Specifically, the encoder network first takes each original training data $x$ from $n_{i,k}^{r}$ and transforms $x$ by mapping $x$ to the network. This process yields the hidden feature representation of $x$, denoted as $z(x)$. The transformation is expressed as $z(x)=q\left(W_{i,e,k}^{r}x+b_{i,e,k}^{r}\right)$, here $W_{i,e,k}^{r}$ and $b_{i,e,k}^{r}$ are the weight matrix and bias vector of the encoder network of UE $i$ in iteration $k$ of round $r$, respectively. To make the paper clearer, we simplify $W_{i,e,k}^{r}$ and $b_{i,e,k}^{r}$ to $W_{e,k}$ and $b_{e,k}$, respectively. We also apply similar simplifications to symbols that will appear later
@@ IV-C1 MADRL Framework
The MADRL framework includes agents, states, actions, rewards and policies[19, 20]. In our cooperative edge caching, each SBS $b$ can be viewed as an agent and makes its action based on the its local state. Then we will define states, actions, rewards and policies at time slot $t$ as follows.
@@ 16             Store transition $(s^{t},a^{t},R^{t},R^{t}_{L},s^{t+1})$ in ()()()()( ) ( );
Next, the algorithm executes for $R_{max}$ episodes. In the first episode, each SBS $b$ obtains $F_{p}$ popular contents $p_{b}^{1}$ from predicted interested contents uploaded by UEs within its own coverage, and randomly selects $C$ contents for caching. Thus each SBS $b$ can obtain the initial caching contents $c_{b}^{1}$. Based on $c_{b}^{1}$ and $p_{b}^{1}$, each SBS $b$ can obtain its initial local state $\bm{s_{b}^{1}}$ according to Eq. (16), and the initial global state can also be further obtained according to Eq. (17), i.e., $\bm{s^{1}}=\left[s_{1}^{1},\ldots,s_{b}^{1},\ldots,s_{B}^{1}\right]$. Then algorithm is iteratively executed from slot $1$ to slot $T$. In the first slot, each SBS $b$ inputs its local state $s_{b}^{1}$ to its local actor network and gets its policy $\pi_{b}^{1}$, then the local action $a_{b}^{1}$ can be obtained based on $\pi_{b}^{1}$. Thus the global action can be obtained according to Eq. (19), i.e., $\bm{a^{1}}=\left[a_{1}^{1},\ldots,a_{b}^{1},\ldots,a_{B}^{1}\right]$. After each SBS $b$ takes the local action $a_{b}^{1}$, it can get its local reward $R_{b}^{1}$ according to Eq. (20). Thus all agents’ rewards can be obtained, i,e., $\bm{R_{L}^{1}}=\left[R_{1}^{1},\ldots,R_{b}^{1},\ldots,R_{B}^{1}\right]$. The global reward $R^{1}$ can also be obtained according to Eq. (21). In addition, each SBS $b$ enters the next local state $s_{b}^{2}$ after taking the local action, and the next global state $s^{2}$ can be further obtained, i.e., $\bm{s^{2}}=\left[s_{1}^{2},\ldots,s_{b}^{2},\ldots,s_{B}^{2}\right]$. Then the transition tuple $\left(\bm{s^{1}},\bm{a^{1}},R^{1},\bm{R_{L}^{1}},\bm{s^{2}}\right)$ is stored in the replay buffer $D$. When the number of transition tuples stored in the replay buffer $D$ is less than $M$, the algorithm starts the next slot.
@@
After that $\varphi_{1}$, $\varphi_{2}$, $\phi_{b}$ and $\theta_{b}$ are updated through gradient ascending based on $\nabla L\left(\varphi_{1}\right)$, $\nabla L\left(\varphi_{2}\right)$, $\nabla L\left(\phi_{b}\right)$ and $\nabla J_{\theta_{b}}$. Then the parameters of the two target global critic networks, local actor network and the critic network are updated as
@@ IV-D Computational Complexity Analysis
Let $G_{E}$, $G_{D}$, $G_{D\textquoteright}$, $G_{A}$, and $G_{C}$ be the computational complexity of computing gradients for the encoder network, the decoder network, the discriminator network, the actor network, and the critic network, respectively. Let $U_{E}$, $U_{D}$, $U_{D\textquoteright}$, $U_{A}$, and $G_{C}$ be the computational complexity of updating parameters for the aforementioned networks. First, we analyze the complexity of the elastic FL training of AAE, which is influenced by the number of SBSs and the number of vehicles covered by each SBS. Let $n_{b}$ represent the number of vehicles under SBS b, and each vehicle needs to undergo $e$ iterations of local training. Therefore, the complexity of the elastic FL training of AAE is $O_{aarfl}=O\left(R_{max}\sum^{B}_{b=1}n_{b}e\left(2G_{E}+G_{D}+G_{D^{\prime}}+2U_{E}+U_{D}+U_{D^{\prime}}\right)\right)$.
@@ V Simulation Results
$R_{max}$
$1000$
$E^{\prime}$
$100$
$M$
$256$
$T$
$100$
$\alpha$
$1$
$\beta$
$30$
$\chi$
$100$
$\delta$
$100$
$\gamma$
$0.99$
$\eta$
$0.01$
$\tau$
$0.001$
@@ V-C Performance Evaluation
For the Random scheme, the complexity is $O(C)$, where $C$ is the cache capacity of the SBS. The C-$\varepsilon$-greedy scheme’s complexity is $O(B\times m\log m)$, with $B$ representing the number of UEs covered by each SBS, and $m$ being the number of content requests by UEs. Thompson sampling has an approximate complexity of $O(e_{t}\times B\times m\log m)$. The BSG scheme’s complexity is approximately $O(m_{all}\log m_{all})$, where $m_{all}$ indicates all cached content in the CS. The complexity of both Traditional Federated MADRL and our CMREF is approximated as $O_{aarfl}+O_{madrl}$. The complexity of EFNRL is approximately $O_{aarfl}$.

### v2

@@ IV-A2 Local Training
The reconstruction stage which mainly focuses on training the AEN is first executed. Specifically, the encoder network first takes each original training data $x$ from $n_{i,k}^{r}$ and transforms $x$ by mapping $x$ to the network. This process yields the hidden feature representation of $x$, denoted as $z(x)$. The transformation is expressed as $z(x)=q\left(W_{i,e,k}^{r}x+b_{i,e,k}^{r}\right)$, here $W_{i,e,k}^{r}$ and $b_{i,e,k}^{r}$ are the weight matrix and bias vector of the encoder network of UE $i$ in iteration $k$ of round $r$, respectively. To make the paper clearer, we simplify $W_{i,e,k}^{r}$ and $b_{i,e,k}^{r}$ to $W_{e,k}$ and $b_{e,k}$, respectively. We also apply similar simplifications to symbols that will appear later.
@@ IV-C1 MADRL Framework
The MADRL framework includes agents, states, actions, rewards and policies[34, 36, 35]. In our cooperative edge caching, each SBS $b$ can be viewed as an agent and makes its action based on the its local state. Then we will define states, actions, rewards and policies at time slot $t$ as follows.
@@ 16             Store transition $(s^{t},a^{t},R^{t},R^{t}_{L},s^{t+1})$ in ()()()()( ) ( );
Next, the algorithm executes for $R_{max}$ episodes. In the first episode, each SBS $b$ obtains $F_{p}$ popular contents $p_{b}^{1}$ from predicted interested contents uploaded by UEs within its own coverage, and randomly selects $C$ contents for caching. Thus each SBS $b$ can obtain the initial caching contents $c_{b}^{1}$. Based on $c_{b}^{1}$ and $p_{b}^{1}$, each SBS $b$ can obtain its initial local state $\bm{s_{b}^{1}}$ according to Eq. (16), and the initial global state can also be further obtained according to Eq. (17), i.e., $\bm{s^{1}}=\left[s_{1}^{1},\ldots,s_{b}^{1},\ldots,s_{B}^{1}\right]$.
Then algorithm is iteratively executed from slot $1$ to slot $T$. In the first slot, each SBS $b$ inputs its local state $s_{b}^{1}$ to its local actor network and gets its policy $\pi_{b}^{1}$, then the local action $a_{b}^{1}$ can be obtained based on $\pi_{b}^{1}$. Thus the global action can be obtained according to Eq. (19), i.e., $\bm{a^{1}}=\left[a_{1}^{1},\ldots,a_{b}^{1},\ldots,a_{B}^{1}\right]$. After each SBS $b$ takes the local action $a_{b}^{1}$, it can get its local reward $R_{b}^{1}$ according to Eq. (20). Thus all agents’ rewards can be obtained, i,e., $\bm{R_{L}^{1}}=\left[R_{1}^{1},\ldots,R_{b}^{1},\ldots,R_{B}^{1}\right]$. The global reward $R^{1}$ can also be obtained according to Eq. (21). In addition, each SBS $b$ enters the next local state $s_{b}^{2}$ after taking the local action, and the next global state $s^{2}$ can be further obtained, i.e., $\bm{s^{2}}=\left[s_{1}^{2},\ldots,s_{b}^{2},\ldots,s_{B}^{2}\right]$.
Then the transition tuple $\left(\bm{s^{1}},\bm{a^{1}},R^{1},\bm{R_{L}^{1}},\bm{s^{2}}\right)$ is stored in the replay buffer $D$. When the number of transition tuples stored in the replay buffer $D$ is less than $M$, the algorithm starts the next slot.
@@
After that $\varphi_{1}$, $\varphi_{2}$, $\phi_{b}$ and $\theta_{b}$ are updated through gradient ascending based on $\nabla L\left(\varphi_{1}\right)$, $\nabla L\left(\varphi_{2}\right)$, $\nabla L\left(\phi_{b}\right)$ and $\nabla J_{\theta_{b}}$.
Then the parameters of the two target global critic networks, local actor network and the critic network are updated as
@@ IV-D Computational Complexity Analysis
Let $G_{E}$, $G_{D}$, $G_{D\textquoteright}$, $G_{A}$, and $G_{C}$ be the computational complexity of computing gradients for the encoder network, the decoder network, the discriminator network, the actor network, and the critic network, respectively. Let $U_{E}$, $U_{D}$, $U_{D\textquoteright}$, $U_{A}$, and $G_{C}$ be the computational complexity of updating parameters for the aforementioned networks. First, we analyze the complexity of the elastic FL training of AAE, which is influenced by the number of SBSs and the number of UEs covered by each SBS. Let $n_{b}$ represent the number of UEs under SBS b, and each UE needs to undergo $e$ iterations of local training. Therefore, the complexity of the elastic FL training of AAE is $O_{aarfl}=O\left(R_{max}\sum^{B}_{b=1}n_{b}e\left(2G_{E}+G_{D}+G_{D^{\prime}}+2U_{E}+U_{D}+U_{D^{\prime}}\right)\right)$.
@@ V Simulation Results
$R_{max}$ $1000$ $E^{\prime}$ $100$
$M$ $256$ $T$ $100$
$\alpha$ $1$ $\beta$ $30$
$\chi$ $100$ $\delta$ $100$
$\gamma$ $0.99$ $\eta$ $0.01$
$\tau$ $0.001$
@@ V-C Performance Evaluation
For the Random scheme, the complexity is $O(C)$, where $C$ is the cache capacity of the SBS. The C-$\varepsilon$-greedy scheme’s complexity is $O(B\times m\log m)$, with $B$ representing the number of UEs covered by each SBS, and $m$ being the number of content requests by UEs. Thompson sampling has an approximate complexity of $O(e_{t}\times B\times m\log m)$. The BSG scheme’s complexity is approximately $O(m_{all}\log m_{all})$, where $m_{all}$ indicates all cached content in the CS. The complexity of both Traditional Federated MADRL and our CEFMR is approximated as $O_{aarfl}+O_{madrl}$. The complexity of EFNRL is approximately $O_{aarfl}$.